please help me solve this is what I got 5x-3y=13
4x-3y=11 My answer is x=3 and y =1 I do not think this is correct. please help.

Answer by funmath(2933)

E1: 5x-3y=13
E2: 4x-3y=11
:
This is the elimination method.
Multiply the E2 by -1 and add it to E1 and the y's will be eliminated so that you can solve for x.
:
-1(4x-3y)=-1(11) --> -4x+3y=-11
:
5x-3y=13
-4x+3y=-11
__________
x+0=2
x=2
:
Substitute 2 in for x in either equation and solve for y.
4(2)-3y=11
8-3y=11
8-8-3y=11-8
-3y=3
-3y/-3=3/-3
y=-1
:
Therefore, your solution is: (x,y)=(2,-1)
Happy Calculating!!!